OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = nine divided by square root of quantity five x plus five. Answers: f(x) = nine divided by square root of x., g(x) = 5x + 5 f(x) = square root of quantity five x plus five., g(x) = 9 f(x) = nine divided by x., g(x) = 5x + 5 f(x) = 9, g(x) = square root of quantity x plus five

OpenStudy (anonymous):

I need helps pleaseeeeee!!!

OpenStudy (anonymous):

What have you attempted thus far?

OpenStudy (anonymous):

nothing, im super lost.

OpenStudy (anonymous):

let f(x)= \[9/\sqrt{x}\] and \[g(x)=5x+5\] so f(g(x)) means that you are plugging g(x) into f's x. so you end up with \[f(g(x))=\frac{ 9 }{ \sqrt{5x+5} }\]

OpenStudy (anonymous):

but whats f(x) & g(x)?

OpenStudy (anonymous):

thats another way to write y but using different letters lets you know that they are two different functions.

OpenStudy (anonymous):

oh okay, I get it now. Thank you!!